OpenStudy (anonymous):

I've tried doing this but I don't get it! Simplify. (–8x) ÷ (–2) A. 2x B. 16x C. 4x D. 03_06_LQ_3d.gif

OpenStudy (johnweldon1993):

\[\large \frac{-8x}{-2}\] Well you know that a negative divided by a negative is positive...so \[\large \frac{8x}{2}\] Now note... \[\large \frac{8x}{2} = \frac{8}{2} \times x = 4x\] I hope that makes sense?

OpenStudy (anonymous):

Sorta makes sense, thank you :)

OpenStudy (johnweldon1993):

When you get down to \[\large \frac{8x}{2}\] all you really do is divide the numbers like normal...but remember at the end...you need to carry along the 'x'... so 8/2 is the 4 ...but we still have the 'x' ...so 4 times x = 4x